### v1

# Stability of Jordan Recurrent Neural Network Estimator

Abstract
State estimation involves finding the states of a noisy dynamical system with a noisy initial condition using noisy measurements, given a model of the system. Many methods exist for state estimation of dynamical systems. For linear dynamical systems with white Gaussian noises of known mean and variance, Kalman filtering is an optimal state estimation algorithm. While several extensions of the Kalman filter to nonlinear systems exist, either the method is not optimal or calculation is not feasible. Several approaches using neural networks have been developed in recent years. While these approaches seem to work numerically, stability and convergence guarantees do not exist. In this paper, we propose to use a Jordan recurrent neural network (JRN) for state estimation. An input-to-state stability (ISS) analysis of the error dynamics is used. Our results are compared to the Kalman filter for the linear case and the extended Kalman filter for the nonlinear case.
I INTRODUCTION
State estimation refers to estimating the states of a noisy dynamical system given noisy and incomplete measurements, partial information about the initial state, as well as a mathematical model for the system. In practical applications, estimating the state of a system is a challenging task for various reasons. The complexity of high-order systems, the availability of a finite number of measurements, inaccurate measurements, incomplete information about the initial state, and errors in the model can make it difficult to estimate the state of the system.
One of the most commonly used optimal state estimation methods is the Kalman filter (KF) [1]. It is optimal for linear systems with white Gaussian noises with zero mean and known covariance. When implemented for a discrete-time system, it consists of two steps: prediction and filtering, and is a recursive algorithm. Many researchers have extended the Kalman filter to nonlinear systems. The most obvious model-based generalization of this filter to nonlinear systems is the Extended Kalman filter (EKF) [2] [3]. The EKF relies on linearization of the dynamical system at each time-step. Although local convergence guarantees [4, 5] exist, convergence to the true state is an issue for highly nonlinear systems and when the estimate of the initial condition is very inaccurate. Moreover, even when the estimation error converges to zero, the method is not optimal.
In recent years, machine learning approaches, particularly artificial neural networks, have attracted attention. Gao et al. [6] used deep long short-term memory (DLSTM) networks to approximate the predicted and filtered state from observations. Although they aimed to study trajectory tracking, their approach turned out to be a state estimation problem since they ignored certain assumptions and used knowledge of the model for training dataset generation. They presented two frameworks, Bayesian and non-Bayesian. For the Bayesian approach, they used Bayes’ rule [7] and designed the DLSTM so that it predicts and filters iteratively within the same network. For the non-Bayesian approach, they considered two DLSTM networks, one for prediction and the other for filtering, based on the fact that neural networks can perform both steps independently. Thus, one network performed prediction without measurements, and the other corrected the predictions by using new measurements. Estimation accuracy, error bounds, analytic performance evaluation, and neural network convergence remain open problems for these approaches. Breitan et al. [8] proposed a machine learning framework that overcame the requirement to find solutions to the Hamilton-Jacobi-Bellman (HJB) partial differential equation for an optimal estimator in the continuous-time case. A neural network was used to approximate the gradient of the solution of the HJB equation. They exploited the HJB equation to construct the observer gain, which coincided with the Kalman-Bucy filter for a linear system. Peralez [9] further proposed a deep learning-based Luenberger estimator design for discrete-time nonlinear systems by approximating the mapping using unsupervised learning. Gençay et al [10] compared the method of least squares, Elman recurrent neural network (ERN), and feedforward neural networks (FNN) for noisy time series data with fully observable states. They concluded that ERNs provided the best estimation properties among the three.
Figure 1: The structure of the Jordan recurrent network proposed for state estimation. It uses output-to-hidden recurrent connections similar to the filter’s dynamical system. The symbols $y^{(t)},a^{(t)},x^{(t)}$ and $\hat{x}^{(t)}$ represent input measurement vector, hidden layer vector, true state vector and estimated state vector respectively at time $t$. Cost function $J^{(t)}$ is the mean squared error at time $t$. The weights and biases of the network are represented by $W_{ay},W_{ax},W_{xa},b_{x}$ and $b_{y}.$
Identifying states entirely on the basis of input-output data is called system identification. Recurrent neural networks have been shown to work well for system identification [11]. In fact, Jordan recurrent networks (JRNs) have been used for system identification by several researchers and have been shown to work as well as ERNs [12][13]. A convergence analysis, with some assumptions, of both networks for system identification has been discussed in [14]. Forecasting is a domain in which future values are predicted based on observed trends. In [15], an Elman-based long short-term memory (LSTM) network is compared with an EKF for time series forecasting. The authors conclude that the LSTM performs better than the EKF for nonlinear systems. In [16], authors show that the ERN performs at least as well as FNN for Lyapunov exponents forecasting, discussing cases where it performs better in detail as well.
This encouraged us to use recurrent neural networks for state estimation. The issue with ERNs is that proving the stability of the estimator is difficult, as the dependence is not directly on the estimated state but on the hidden state at the previous time-step. This is not the case for a JRN. We use an algorithm that uses known information about a dynamical system to train a JRN for state estimation. The structure of JRNs resembles that of a dynamical system, making it a suitable candidate for state estimation. The difference between Elman and Jordan recurrent neural networks lies in the types of recurrent connections [17]. Schäfer and Zimmerman [18] proved ERNs are universal approximators of dynamical systems. These results can be extended to JRNs for state estimation [19].
The stability of recurrent neural networks (RNNs) in the sense of Lyapunov has been well-developed in [20, 21], although the stability of errors in RNNs as neural estimators for state estimation has not been extensively explored. Finding Lyapunov functions is generally challenging. Several papers [22, 23] have shown that Lyapunov functions can be attained and represented by expressions of the compositional structure of the neural network, the correctness being guaranteed by satisfiability modulo theories (SMT) solvers. In this paper, we design an estimator using JRNs and employ an input-to-state stability (ISS) approach to analyze the stability of the error dynamics. We provide a theoretical result establishing a condition under which the error dynamics will converge to zero if the original system is stable. Estimation of unstable systems is generally conducted in combination with a stabilizing feedback, so our approach assumes that any necessary stabilization has been performed. The corresponding Lyapunov functions are learned by FNNs using an ISS approach. To the best of the authors’ knowledge, this is the first work analyzing the stability of errors using ISS and finding the ISS Lyapunov function with a neural network for state estimation.
We use a hybrid approach that combines both model and data. We employ a shallow JRN, presented in Section II and Section III-A, to estimate the states of the dynamical system. Data for training the recurrent network is model-based and generated as explained in Section II-B. This ensures that the training data reflect the assumptions of a Kalman filter and an extended Kalman filter. Input-to-state stability analysis of the Jordan recurrent network-based estimator is discussed in Section II, while the method to find the corresponding Lyapunov functions is explained in Section II-B. The approach is then tested in Section IV on three different discrete-time systems, obtained from the discretization of ODEs, one linear and two nonlinear. This is followed by a discussion of the convergence and input-to-state stability of the results.
II Input-to-state stability analysis
Consider a discrete-time nonlinear system
$\begin{split}x^{(t+1)}&=f(x^{(t)})+w^{(t+1)},\\ y^{(t+1)}&=h(x^{(t+1)})+v^{(t+1)},\\ x^{(0)}&=x^{0}+\bar{x}^{0},\end{split}$ (1)
where $x^{(t)}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state vector at time-step $t$, $w^{(t+1)}\in\mathbb{R}^{n}$ is the process noise vector at time-step $t+1$, $y^{(t+1)}\in\mathcal{Y}\subseteq\mathbb{R}^{m}$ is the measurement vector at time-step $t+1$, $v^{(t+1)}\in\mathbb{R}^{m}$ is the measurement noise vector at time-step $t+1$, $x^{0}\in\mathbb{R}^{n}$ is the true initial condition and $\bar{x}^{0}\in\mathbb{R}^{n}$ is the initial condition noise. Here, $\mathcal{X}$ and $\mathcal{Y}$ are the state space and measurement space, respectively, and $n\in\mathbb{N}$ is the total number of states and $m\in\mathbb{N}$ is the number of states whose noisy measurements are available. The functions $f$ and $h$ are assumed to be sufficiently smooth.
Our aim is to estimate the state vector $x^{(t+1)}$ by $\hat{x}^{(t+1)}$ based on the measurement vector $y^{(t+1)},$ previous state vector estimate $\hat{x}^{(t)}$ and the unknown noisy initial condition $x^{(0)}$.
A JRN has recurrent connections from the output of the previous layer to the hidden state of the next layer. Thus, we use $y^{(t+1)}$ and $\hat{x}^{(t+1)}$ as inputs and outputs, respectively, of the JRN. The forward propagation of the network is
$\begin{split}a^{(t+1)}&=\sigma(W_{ay}y^{(t+1)}+W_{ax}\hat{x}^{(t)}),\\ \hat{x}^{(t+1)}&=W_{xa}a^{(t+1)},\end{split}$ (2)
where $\sigma$ is the activation function and $\hat{x}^{(t+1)}$ is the estimated state vector at time-step $t+1$. The hidden layer vector at time-step $t+1$ is given by $a^{(t+1)}.$ The weights and biases of the network are represented by $W_{ay},W_{ax}$ and $W_{xa}$. The network is depicted in Fig. 1. This is different from an ERN because the recurrent connections for an ERN are from the previous hidden layer $a^{(t)}$ to the next hidden layer $a^{(t+1)}$, whereas for a JRN, they are from the previous output $\hat{x}^{(t)}$ to next hidden layer $a^{(t+1)}.$ Thus, for an ERN, the term $W_{ax}\hat{x}^{(t)}$ would be replaced by $W_{aa}{a}^{(t)}$ where $W_{aa}$ is a weight matrix. Note that we are considering a bias-free network.
To analyze the stability of the neural estimator, we first consider the case where there are no process noise and measurement noise in (1), i.e., $w^{(t+1)}=v^{(t+1)}=0$ for all $t$, and $\bar{x}^{0}=0$, as follows,
$\begin{split}x^{(t+1)}&=f(x^{(t)}),\\ y^{(t+1)}&=h(x^{(t+1)}).\end{split}$ (3)
The solution of system (3) is denoted by $x^{(t)}(\xi)$ at time step $t$ with initial condition $x^{(0)}=\xi\in\mathcal{X}$. We define the error term as $e^{(t+1)}=x^{(t+1)}-\hat{x}^{(t+1)}$. Then,
$\begin{split}e^{(t+1)}&=x^{(t+1)}-W_{xa}\sigma(W_{ay}y^{(t+1)}+W_{ax}\hat{x}^{(t)})\\ &=f(x^{(t)})-W_{xa}\sigma(W_{ay}h(f(x^{(t)}))\\ &\quad+W_{ax}x^{(t)}-W_{ax}e^{(t)}).\end{split}$ (4)
Obviously, $e^{(t+1)}$ is a function of $e^{(t)}$ and $x^{(t)}$, and we call it the error system, denoted as follows,
$e^{(t+1)}\vcentcolon=g(e^{(t)},x^{(t)}),$ (5)
where $e^{(t)}\in\mathcal{E}\subseteq\mathbb{R}^{n}$, and $\mathcal{E}$ is the space for the error term. In a similar manner, we denote the solution to system (5) as $e^{(t)}(\eta,x)$ with the initial condition $e^{(0)}=\eta\in\mathcal{E}$.
We introduce Lyapunov functions for the error dynamics, regarding $x^{(t)}$ as the input in (5).
Definition II.1
A continuous function $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{K}$-function if it is strictly increasing and $\alpha(0)=0$. It is a $\mathcal{K}_{\infty}$-function if it is a $\mathcal{K}$-function and $\alpha(r)\rightarrow\infty$ as $r\rightarrow\infty.$
Definition II.2
A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{KL}$-function if, for each fixed $t$, the function $\beta(\cdot,t)$ is a $\mathcal{K}$-function with respect to $r$, and for each fixed $r$, the function $\beta(r,\cdot)$ is decreasing with respect to $t$, and $\beta(r,t)\rightarrow 0$ as $t\rightarrow\infty.$
Definition II.3
The origin of (3) is globally asymptotically stable if, for each $\varepsilon>0$, there is a $\delta=\delta(\varepsilon)>0$ such that
$\|\xi\|<\delta\implies\|x^{(t)}\|<\varepsilon,\forall t\geq 0,\text{and}\lim_{t\rightarrow\infty}x^{(t)}=0.$ (6)
In addition, there exists a $\mathcal{KL}$-function $\beta$ such that
$\|x^{(t)}(\xi)\|\leq\beta(\|\xi\|,t)\quad\forall\xi\in\mathcal{X},\forall t\in\mathbf{Z}_{+}.$ (7)
Definition II.4
[24] The discrete-time system (5) is input-to-state stable (ISS) if there exists a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\gamma$, such that for each $t\in\mathbf{Z}_{+}$, the following inequality holds,
$|e^{(t)}(\eta,x)|\leq\beta(\|\eta\|,t)+\gamma(\|x\|),$ (8)
for all $\eta\in\mathcal{E}$ and for all $x\in\mathcal{X}$.
Definition II.5
[24] A continuous function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$ is called an ISS-Lyapunov function for the discrete-time system (5) if there exist $\mathcal{K}_{\infty}$-functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a $\mathcal{K}$-function $\gamma$ such that
$\alpha_{1}(\|e^{(t)}\|)\leq V(e^{(t)})\leq\alpha_{2}(\|e^{(t)}\|),\quad\forall e^{(t)}\in\mathcal{E},$ (9)
and
$\begin{split}V(g(e^{(t)},x^{(t)}))-V(e^{(t)})&\leq-\alpha_{3}(\|e^{(t)}\|)+\gamma(\|x^{(t)}\|)\\ &\forall e^{(t)}\in\mathcal{E},\forall x^{(t)}\in\mathcal{X}.\end{split}$ (10)
Theorem II.6
If the error system (5) with $x^{(t)}$ as input is ISS and the origin of the discrete-time system (3) is globally asymptotically stable, then the origin of the cascade system (3) and (5) is globally asymptotically stable.
Proof:
The solutions of (3) and (5) satisfy:
$\|x^{(t)}(\xi)\|\leq\beta_{1}(\|\xi\|,t),$ (11)
$\|e^{(t)}(\eta,x)\|\leq\beta_{2}(\|\eta\|,t)+\gamma(\|x^{(t)}\|),$ (12)
where $\beta_{1},\beta_{2}$ are $\mathcal{KL}$-functions. Thus,
$\|e^{(t)}(\eta,x)\|\leq\beta_{2}(\|\eta\|,t)+\gamma(\beta_{1}(\|\xi\|,t)).$ (13)
Let $s^{(t)}$ denote the concatenation of the state $x^{(t)}$ and the error state $e^{(t)}$, and $\zeta$ denote the origin of this cascade system. We have $\|x^{(t)}\|\leq\|s^{(t)}\|$ and $\|e^{(t)}\|\leq\|s^{(t)}\|$, and $\|s^{(t)}\|\leq\|x^{(t)}\|+\|e^{(t)}\|$. Defining $\beta(\cdot,\cdot)=\beta_{1}(\cdot,\cdot)+\beta_{2}(\cdot,\cdot)+\gamma(\beta_{1}(\cdot,\cdot))$, this yields
$\|s^{(t)}(\zeta)\|\leq\beta(\|\zeta\|,t)\quad\forall t\in\mathbf{Z}_{+}.$ (14)
It can be easily verified that $\beta$ is a $\mathcal{KL}$-function, which completes the proof. ∎
Two different approaches were used for finding the Lyapunov function, depending on whether the system is linear or nonlinear.
II-A Linear systems
Consider a linear discrete-time system defined by, for matrices $A$ and $H$,
$\begin{split}x^{(t+1)}&=Ax^{(t)},\\ y^{(t+1)}&=Hx^{(t+1)}.\end{split}$ (15)
By using the identity function as the activation function $\sigma$ in (2), the error dynamics of the linear state estimator can be written as
$\begin{split}e^{(t+1)}&=Ax^{(t)}-W_{xa}(W_{ay}y^{(t+1)}+W_{ax}\hat{x}^{(t)})\\ &=Ax^{(t)}-W_{xa}(W_{ay}Hx^{(t+1)}+W_{ax}(x^{(t)}-e^{(t)}))\\ &=(A-W_{xa}W_{ay}HA-W_{xa}W_{ax})x^{(t)}+W_{xa}W_{ax}e^{(t)}.\end{split}$ (16)
Defining $\mathcal{A}=W_{xa}W_{ax}$ and $\mathcal{B}=A-W_{xa}W_{ay}HA-W_{xa}W_{ax}$, the error system is a linear system with $x^{(t)}$ as the input:
$e^{(t+1)}=\mathcal{A}e^{(t)}+\mathcal{B}x^{(t)}.$ (17)
If this discrete-time system is ISS then there is a quadratic Lyapunov function
$V(e)=e^{T}Pe,$ (18)
where $P$ is a positive definite matrix obtained by solving
$\mathcal{A}^{T}P\mathcal{A}-\mathcal{A}+Q=0.$ (19)
Here, $Q$ is a symmetric positive definite matrix. In this case, it is easy to show that both properties in Def. II.5 are satisfied with $\alpha_{1}(x)=\lambda_{min}(P)x^{2}$, $\alpha_{2}(x)=\lambda_{max}(P)x^{2}$, $\alpha_{3}(x)=\frac{1}{2}\lambda_{min}(Q)x^{2}$, and $\gamma(x)=(\frac{2|\mathcal{A}^{T}P\mathcal{B}|^{2}}{\lambda_{min}(Q)}+|\mathcal{B}^{T}P\mathcal{B}|^{2})x^{2}$  [24].
II-B Nonlinear systems
For nonlinear systems, a counterexample-guided method with verification provided by SMT solvers was used to synthesize the ISS-Lyapunov functions for the error system (5).
Inspired by [25], we use a one-hidden layer feed-forward neural network with zero bias terms for all layers to learn a Lyapunov function of the following form
$V(e;\theta)=\sigma(W_{2}\sigma(W_{1}e)),$ (20)
where $W_{1}$ and $W_{2}$ are the weights for the hidden layer and output layer respectively, $\theta$ denotes all the hyper-parameters, $\sigma$ is the activation function, $e$ is the input vector, which is the shorthand notation for the error state at time $t$. Similarly, $x$ denotes the state at time $t$.
We use the square function as the activation function $\sigma$ in this network, which results in sum-of-squares (SOS)-like quadratic Lyapunov functions. Consequently, we have $V(0)=0$. As a result, a valid ISS Lyapunov function is attained when properties (9) and  (10) are satisfied. Its falsification constraints can then be written as a first-order logic formula over the real numbers. This yields
$\begin{split}&\left(\sum_{i=1}^{n}e_{i}^{2}\leq r_{e}\vee\sum_{i=1}^{n}x_{i}^{2}\leq r_{x}\right)\wedge\\ &\bigg{(}(V(x)-\alpha_{1}(|e|)\leq 0)\vee(V(x)-\alpha_{2}(|e|)\geq 0)\vee\\ &(V(g(e,x))-V(e)+\alpha_{3}(|e|)-\gamma(|x|)\geq 0)\bigg{)},\end{split}$ (21)
where, $r_{e}$ and $r_{x}$ are the radii of the so-called valid region for the error state and the state respectively, on which we verify the condition using SMT solvers. When the SMT solver returns UNSAT, a valid ISS Lyapunov function is obtained. Otherwise, it produces counterexamples that can be added to the training dataset of the neural network for finding Lyapunov function candidates. It is worth mentioning that due to the nature of the SMT solvers, we typically need to exclude a small region around the origin when verifying the falsification constraints. Around the origin, the linearization of the nonlinear system dominates, and we can implement the method for the linearized model to provide stability guarantees as detailed in Section II-A. We refer the readers to [26] for a detailed rigorous proof and algorithm.
The loss function is consistent with the falsification constraints:
$\begin{split}L(\theta)&=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\max\bigg{(}0,V_{\theta}(g(e_{i},x_{j}))-V_{\theta}(e_{i})\\ &+\alpha_{3}(|e_{i}|)-\gamma(|x_{j}|)\bigg{)}+\max\bigg{(}0,V_{\theta}(e_{i})-\alpha_{2}(|e_{i}|)\bigg{)}\\ &+\max\bigg{(}0,\alpha_{1}(|e_{i}|)-V_{\theta}(e_{i})\bigg{)}.\end{split}$ (22)
This is known as the positive penalty for the violation of conditions in Definition II.5.
III Implementation
The initial conditions are sampled over an interval. The initial condition is considered to be Gaussian. The covariance of the initial condition is $(0.01)\times I$ where $I$ is an identity matrix of appropriate dimensions. Similarly, process and measurement noises are assumed to be Gaussian with zero mean and covariance $0.01\times I$ where matrix $I$ is of appropriate dimensions. This is done for each sequence in the data set. Since the approach is state estimation, knowledge of the functions $f$ and $h$ in (1) is used in data generation.
The data set is divided into three parts: training, validation and testing in the ratio $80:10:10$
Figure 2: The graph represents a comparison between JRN and KF Error$(t),0\leq t\leq 100$ values of Jordan recurrent network based state estimation and the Kalman filter based state estimation for mass-spring damper system. The magnified graph within the graph provides a closer look at the error values from time $19$ to $20$ seconds. The x-axis of both the original and the magnified graphs represents time in seconds while the y-axis represents error at time $t.$ The blue dotted line graph represents the error values for JRN based estimates while the orange solid line represents the error values for the Kalman filter based estimates.
III-A Training and validation
A custom network with forward propagation as in (2) is implemented. The weight matrix $W_{ay}$ is initialized using Xavier uniform distribution while weight matrices $W_{ya}$ and $W_{xa}$ are initialized to be (semi) orthogonal matrices using Pytorch’s $torch.nn.init$ module. The function $\sigma(z)$ is chosen to be equal to $z$ for linear systems and is chosen to be $\tanh(z)$ for nonlinear systems. The backward propagation is implemented using PyTorch’s $backward()$ function. The loss function at time-step $t+1$ for each sequence is
$J_{i}^{(t+1)}(\phi)=\frac{1}{n}\sum_{i=1}^{n}{(x^{(t+1)}_{i}-\hat{x}^{(t+1)}_{i,\phi})}^{2},$ (23)
where $1\leq i\leq n$ and $n$ is the number of states, $x^{(t+1)}_{i}$ represents the true value of state $x_{i}$ at time-step $t+1$ and $\hat{x}^{(t+1)}_{i,\phi}$ represents the estimated value of state $x_{i}$ at time-step $t+1$ by the JRN where $\phi$ denotes hyper-parameters of the network. Adam optimization is used for training the network. The optimal learning rate is chosen based on the network’s performance over a range of learning rates. The number of hidden units considered is $50$ and a batch size of $40$ is used for each example in Section IV.
Early stopping with a fixed patience value to decide the number of epochs needed is used. In each epoch, network’s performance on validation data is used as the measure to decide whether to proceed or keep training for another epoch.
III-B Testing and comparison with KF and EKF
We tested our estimators for 3 different systems in Section IV by comparing the results with KF for the linear system and EKF for the nonlinear systems. Standard implementations of KF and EKF are used, see for example [2].
As mentioned earlier, $10\%$ of the data set is used for testing. We compare JRN and EKF graphically using average error at time $t$ over all features and all test sequences
$\text{Error}(t)=\frac{1}{m_{test}n}\sum_{k=1}^{m_{test}}\sum_{i=1}^{n}(x_{i}^{(t)[k]}-\hat{x}_{i}^{(t)[k]})^{2},$ (24)
where $m_{test}$ is the number of test sequences, $n$ is the number of states, $x_{i}^{(t)[k]}$ represents the true value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(t)[k]}$ represents the estimated value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence. We also use normalised mean square error (NMSE) to compare different methods. It is calculated as
$\text{NMSE}=\frac{1}{m_{test}nT}\sum_{k=1}^{m_{test}}\sum_{j=1}^{T}\sum_{i=1}^{n}(x_{i}^{(j)[k]}-\hat{x}_{i}^{(j)[k]})^{2},$ (25)
where $m_{test}$ is the number of test sequences, $T$ is the number of time-steps in each sequence, $n$ is the number of states, $x_{i}^{(j)[k]}$ represents the true value of state $x_{i}$ at time-step $j$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(j)[k]}$ represents the estimated value of state $x_{i}$ at time-step $j$ for the $k^{th}$ sequence.
IV Numerical results
In this section, we illustrate the proposed method with several examples. All three discrete-time systems were obtained by discretizing common ordinary differential equation examples. We use a zero order hold discretization for linear systems. For nonlinear continuous-time state space systems, we use the RK-45 discretization using Python’s $scipy.integrate.solve\_ivp$ function.
A remote Windows server with T4 GPU, courtesy of math faculty computing facility (MFCF), University of Waterloo, was used for all numerical simulations.
(a) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 50$ seconds for down pendulum.
(b) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 20$ seconds for reversed Van der Pol oscillator.
Figure 3: The graphs represent a comparison of Error$(t)$ values of Jordan recurrent network-based state estimation and extended Kalman filter-based state estimation for nonlinear dynamical systems. The magnified graph within each graph provides a closer look at the error values from time $19$ to $20$ seconds. The x-axis of both the original and the magnified graphs represents time in seconds, while the y-axis represents error at time $t$. The blue dotted line graph represents the error values for JRN-based estimates, while the orange solid line represents the error values for extended Kalman filter-based estimates.
IV-A Mass-Spring-Damper system
Consider a zero order hold discretization with $\Delta t=0.1s$ of the mass-spring damper system
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{k}{m}x_{1}(t)-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (26)
where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity respectively of the spring. The parameters are $m=10kg,$ $b=6kg/sec$ and $k=800kg/{sec}^{2}.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $100$ sequences from 00 to $100$ seconds each. The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ Early stopping was implemented with a patience of $5$. We use Adam optimization for training with a learning rate of $10^{-4}.$ There were $4216$ epochs for training.
Fig. 2 shows the error at each time-step averaged all test sequences for JRN and EKF. As can be seen from Fig. 2, the error for JRN converges to 00 as fast as for the KF. Also, the NMSE for JRN is $0.0631$ and for KF is $0.0653$ which is very close. Recall that KF is the optimal estimator for this system. Since the initial condition is not assumed to be known by the estimator, there is a spike in the initial values of the error for the KF and the JRN but the spike in KF error is considerably larger than that of JRN. The magnified graph in Fig. 2 shows that after convergence JRN error is very close to that the KF error.
For this 2-dimensional linear system, the origin is asymptotically stable. We then implement the method in II-A for linear systems. Setting $Q=I$, it is easy to establish that $P=\begin{bmatrix}1.0236&-0.0934\\ -0.0934&69.4428\end{bmatrix}$ With this ISS-Lyapunov function, by Theorem II.6, both the original system and the error system are asymptotically stable. This proves the stability of the NN estimator error.
IV-B Down pendulum
Consider an RK-45 discretization with $\Delta t=0.01s$ of the down pendulum
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{g}{l}\sin({x_{1}(t)})-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (27)
where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity of the pendulum respectively. The parameters are $g=9.8m/{sec}^{2},$ $m=2kg,$ $b=0.9kg/sec$ and $l=1m.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $200$ sequences from 00 to $50$ seconds each. The mean of the initial condition is chosen randomly from $[-2,2]\times[-2,2].$ Early stopping was implemented with a patience of $50$. The validation loss reached early stopping patience at $2984$ epochs. We further used Adam optimization for training with a learning rate of $10^{-4}.$
The error at each time-step, averaged over all features and all test sequences, is shown for JRN and EKF in Fig. 3(a). The error for JRN converges to 00 as fast as that for the EKF. Also, the NMSE for JRN is $0.0627$ while for EKF it is $0.0647$ The magnified graph in Fig. 3(a) shows that after convergence JRN error and EKF error are quite close. The initial spike in JRN error is not as high as the EKF error. The error becomes very small as $t$ goes to $\infty$, despite the presence of noise.
For this nonlinear system, an ISS Lyapunov function was learned using a neural network, as shown in Fig. 4, which is verified by an SMT solver, dReal [27], on a $[-4,4]$ for the error state. In this case, we used $\alpha_{1}(\cdot)=\alpha_{3}(\cdot)=0.01|\cdot|$ and $\alpha_{2}(\cdot)=\gamma(\cdot)=100|\cdot|$.
Figure 4: The learned ISS Lyapunov function for the down pendulum example.
IV-C Reversed Van der Pol oscillator
Consider an RK-45 discretization with $\Delta t=0.1s$ of the reversed Van der Pol oscillator
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}-x_{2}(t)\\ x_{1}(t)+(({x_{1}}(t))^{2}-1)x_{2}(t)\end{bmatrix}$ (28)
where $x_{1}(t)$ and $x_{2}(t)$ is the position and velocity of the oscillator. The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $200$ sequences from 00 to $20$ seconds each. The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ The training was done for $35000$ epochs as early stopping patience of $7000$ was not achieved. The validation error continued to decrease. Adam optimization is used for training with a learning rate of $10^{-4}.$ As can be seen from Fig. 3(b), the JRN error converges to zero much faster than the EKF error. Also, the NMSE for JRN is $0.0484$ while for EKF is $0.1025$ This is considerable difference. The magnified graph shows that even after convergence, the JRN error is better than EKF error. This example suggests that when a network is trained for a larger number of epochs, the JRN error can converge much faster than the EKF error.
For this system, in a similar manner to the previous example, we synthesized an ISS-Lyapunov function on $[-2,2]$ for the error states, which provides the asymptotic stability for the error term by Theorem II.6.
V Conclusions and future work
In this paper, we demonstrated that when knowledge of functions $f$ and $h$ in (1) is available or the data of the real systems can be properly sampled, state estimation can be performed by using Jordan recurrent networks. The Jordan recurrent network based estimator helped us achieve results as good as KF and EKF. We also showed the stability of the error dynamics of the neural estimator using an ISS approach, by taking advantage of the structural properties of an unbiased JRN. For linear systems, quadratic Lyapunov equations can be computed analytically for the error dynamics to provide stability. For nonlinear systems, a counter-example guided method was used to learn an ISS Lyapunov function. Furthermore, the effectiveness of the proposed method has been illustrated with three 2-dimensional systems. We will extend our approach to higher dimensional systems in future research. Another potential future research direction is to consider stability for the case where biases are non-zero. Additional further work is to extend the results on the convergence of discrete-time recurrent neural networks as shown in [28] to JRNs.

### v2

@@
 Abstract
-State estimation involves finding the states of a noisy dynamical system with a noisy initial condition using noisy measurements, given a model of the system.
-Many methods exist for state estimation of dynamical systems.
-For linear dynamical systems with white Gaussian noises of known mean and variance, Kalman filtering is an optimal state estimation algorithm.
-While several extensions of the Kalman filter to nonlinear systems exist, either the method is not optimal or calculation is not feasible.
-Several approaches using neural networks have been developed in recent years.
-While these approaches seem to work numerically, stability and convergence guarantees do not exist.
-In this paper, we propose to use a Jordan recurrent neural network (JRN) for state estimation.
-An input-to-state stability (ISS) analysis of the error dynamics is used.
-Our results are compared to the Kalman filter for the linear case and the extended Kalman filter for the nonlinear case.
+State estimation refers to determining the states of a dynamical system that starts from a noisy initial condition and evolves under process noise, based on noisy measurements and a known system model.
+For linear dynamical systems with white Gaussian noises of known mean and variance, Kalman filtering is a well-known method that leads to stable error dynamics for detectable systems.
+There are some non-optimal extensions to nonlinear systems.
+Recent work has used neural networks to develop estimators for nonlinear systems that optimize a criterion.
+Stability of the error dynamics is even more important than optimality.
+Jordan recurrent neural networks (JRNs) have a structure that mimics that of a dynamical system and are thus appealing for estimator design.
+We show that a JRN performs better than an extended Kalman filter(EKF) and unscented Kalman filter(UKF) for several examples.
+The main contribution of this paper is an input-to-state stability analysis of the error dynamics of JRNs.
+The stability of the error dynamics of several examples is shown.
 I INTRODUCTION
-State estimation refers to estimating the states of a noisy dynamical system given noisy and incomplete measurements, partial information about the initial state, as well as a mathematical model for the system.
-In practical applications, estimating the state of a system is a challenging task for various reasons.
-The complexity of high-order systems, the availability of a finite number of measurements, inaccurate measurements, incomplete information about the initial state, and errors in the model can make it difficult to estimate the state of the system.
-One of the most commonly used optimal state estimation methods is the Kalman filter (KF) [1].
-It is optimal for linear systems with white Gaussian noises with zero mean and known covariance.
-When implemented for a discrete-time system, it consists of two steps: prediction and filtering, and is a recursive algorithm.
-Many researchers have extended the Kalman filter to nonlinear systems.
-The most obvious model-based generalization of this filter to nonlinear systems is the Extended Kalman filter (EKF) [2] [3].
-The EKF relies on linearization of the dynamical system at each time-step.
-Although local convergence guarantees [4, 5] exist, convergence to the true state is an issue for highly nonlinear systems and when the estimate of the initial condition is very inaccurate.
-Moreover, even when the estimation error converges to zero, the method is not optimal.
-In recent years, machine learning approaches, particularly artificial neural networks, have attracted attention.
-Gao et al.
-[6] used deep long short-term memory (DLSTM) networks to approximate the predicted and filtered state from observations.
-Although they aimed to study trajectory tracking, their approach turned out to be a state estimation problem since they ignored certain assumptions and used knowledge of the model for training dataset generation.
-They presented two frameworks, Bayesian and non-Bayesian.
-For the Bayesian approach, they used Bayes’ rule [7] and designed the DLSTM so that it predicts and filters iteratively within the same network.
-For the non-Bayesian approach, they considered two DLSTM networks, one for prediction and the other for filtering, based on the fact that neural networks can perform both steps independently.
-Thus, one network performed prediction without measurements, and the other corrected the predictions by using new measurements.
-Estimation accuracy, error bounds, analytic performance evaluation, and neural network convergence remain open problems for these approaches.
-Breitan et al.
-[8] proposed a machine learning framework that overcame the requirement to find solutions to the Hamilton-Jacobi-Bellman (HJB) partial differential equation for an optimal estimator in the continuous-time case.
-A neural network was used to approximate the gradient of the solution of the HJB equation.
-They exploited the HJB equation to construct the observer gain, which coincided with the Kalman-Bucy filter for a linear system.
-Peralez [9] further proposed a deep learning-based Luenberger estimator design for discrete-time nonlinear systems by approximating the mapping using unsupervised learning.
-Gençay et al [10] compared the method of least squares, Elman recurrent neural network (ERN), and feedforward neural networks (FNN) for noisy time series data with fully observable states.
-They concluded that ERNs provided the best estimation properties among the three.
-Figure 1: The structure of the Jordan recurrent network proposed for state estimation.
-It uses output-to-hidden recurrent connections similar to the filter’s dynamical system.
-The symbols $y^{(t)},a^{(t)},x^{(t)}$ and $\hat{x}^{(t)}$ represent input measurement vector, hidden layer vector, true state vector and estimated state vector respectively at time $t$.
-Cost function $J^{(t)}$ is the mean squared error at time $t$.
-The weights and biases of the network are represented by $W_{ay},W_{ax},W_{xa},b_{x}$ and $b_{y}.$
-Identifying states entirely on the basis of input-output data is called system identification.
+State estimation refers to estimating the states of a dynamical system given a model and some measurements; both of which are subject to errors and disturbances.
+It differs from system identification where there is no mathematical model.
+This is needed in some situations.
+However, often a model for the dynamics, although not perfect is available, and should be used in estimator design.
+The need for an estimator to run in real-time despite complex dynamics, inaccurate measurements, incomplete information about the initial state, and errors in the model all mean that estimation, particularly of nonlinear systems, is challenging.
+Several approaches to state estimation exist dating from the early 1960s.
+Kalman filtering (KF) is a well known and widely used method for linear systems.
+If the disturbances on the model and the measurements are independent Gaussian noises, with known means and covariances then the Kalman filter yields an estimate of minimum error covariance.
+For nonlinear systems, several extensions of the Kalman filter are commonly used.
+In an extended Kalman filter (EKF), the system is linearised at each time-step and a Kalman filter for the linearized system is used [1, e.g.].
+It is widely used due to its simplicity and low computational overhead.
+Another extension is the unscented Kalman filter (UKF).
+This method relies on choosing sigma points to propagate state mean and covariance [1, e.g.].
+The UKF typically copes better with highly nonlinear systems, but is computationally expensive for high-order systems.
+Machine learning approaches are increasingly common.
+A deep long short-term memory network has been used to approximate the predicted and filtered state using observations in [2].
+As for optimal control, a Hamilton-Jacobi-Bellman partial differential equation needs to be solved to find the minimum variance (or maximum likelihood) estimator.
+Its solution is very difficult for nonlinear systems since the number of independent variables equals the order of the system (the curse of dimensionality).
+In [3], a neural network is used to find solutions to the Hamilton-Jacobi-Bellman equation, assuming that a smooth solution exists.
+In [4], a non-optimal design for discrete-time nonlinear systems is implemented using unsupervised learning.
+In another work, [5] compared the method of least squares, Elman recurrent neural network (ERN) and feedforward neural network (FNN) for noisy time series data with fully observable states and concluded that ERN provides the best estimation.
+Other papers that explore RNNs for state estimation include [6, 7, 8, 9].
 Recurrent neural networks have been shown to work well for system identification[10].
-In fact, Jordan recurrent networks (JRNs) have been used for system identification by several researchers and have been shown to work as well as ERNs [12][13].
-A convergence analysis, with some assumptions, of both networks for system identification has been discussed in [14].
-Forecasting is a domain in which future values are predicted based on observed trends.
-In [15], an Elman-based long short-term memory (LSTM) network is compared with an EKF for time series forecasting.
-The authors conclude that the LSTM performs better than the EKF for nonlinear systems.
-In [16], authors show that the ERN performs at least as well as FNN for Lyapunov exponents forecasting, discussing cases where it performs better in detail as well.
-This encouraged us to use recurrent neural networks for state estimation.
-The issue with ERNs is that proving the stability of the estimator is difficult, as the dependence is not directly on the estimated state but on the hidden state at the previous time-step.
+Jordan recurrent networks (JRNs) have been used for system identification by several researchers and have been shown to work as well as ERNs [11, 12, 13].
+In [14] the structure of a Jordan recurrent network was extended to an LSTM and compared to an Elman long short-term memory network for estimator design.
+The training time considerably decreased when using JLSTM for state estimation.
+This is consistent with the fact that the structure of a Jordan RNN is closer to that of a dynamical system than that of a Elman RNN [15].
+Just as stability is critical for a control system, in the absence of disturbances, the estimated state must converge to the true state.
+Small disturbances should increase the error by a small amount.
+In other words, the error dynamics must be input-to-state stable (ISS) [16, 17].
+It is known that for a linear system, if the measurements satisfy the weak assumption of detectability, the error dynamics are stable.
+It was recently shown [18] that if a nonlinear system is locally detectable, then the error dynamics of an EKF are locally ISS.
+If the system is uniformly observable and can be put globally into the normal form, along with some other technical assumptions on the system dynamics then the EKF error dynamics are stable [19, sec.
+2.4].
+For a discrete-time UKF satisfying certain assumptions, the estimation error remains bounded [20].
+The stability of errors of RNNs as state estimators has not been well explored.
+However, there are several approaches to establishing the stability of RNNs by finding a Lyapunov function [21, 22, 23].
+This suggested extending this work to show stability of error dynamics.
+The issue with ERNs in this context is that proving the stability of the estimator is difficult due to its dependence on the hidden neuron state at the previous time-step.
 This is not the case for a JRN.
-We use an algorithm that uses known information about a dynamical system to train a JRN for state estimation.
-The structure of JRNs resembles that of a dynamical system, making it a suitable candidate for state estimation.
-The difference between Elman and Jordan recurrent neural networks lies in the types of recurrent connections [17].
-Schäfer and Zimmerman [18] proved ERNs are universal approximators of dynamical systems.
-These results can be extended to JRNs for state estimation [19].
-The stability of recurrent neural networks (RNNs) in the sense of Lyapunov has been well-developed in [20, 21], although the stability of errors in RNNs as neural estimators for state estimation has not been extensively explored.
+We therefore design an estimator using JRNs and utilize an input-to-state stability (ISS) approach to analyze the stability of the error dynamics.
 Finding Lyapunov functions is generally challenging.
 Several papers [24, 23] have shown that Lyapunov functions can be attained and represented by expressions of the compositional structure of the neural network, the correctness being guaranteed by satisfiability modulo theories (SMT) solvers.
-In this paper, we design an estimator using JRNs and employ an input-to-state stability (ISS) approach to analyze the stability of the error dynamics.
-We provide a theoretical result establishing a condition under which the error dynamics will converge to zero if the original system is stable.
-Estimation of unstable systems is generally conducted in combination with a stabilizing feedback, so our approach assumes that any necessary stabilization has been performed.
-The corresponding Lyapunov functions are learned by FNNs using an ISS approach.
-To the best of the authors’ knowledge, this is the first work analyzing the stability of errors using ISS and finding the ISS Lyapunov function with a neural network for state estimation.
-We use a hybrid approach that combines both model and data.
-We employ a shallow JRN, presented in Section II and Section III-A, to estimate the states of the dynamical system.
-Data for training the recurrent network is model-based and generated as explained in Section II-B.
-This ensures that the training data reflect the assumptions of a Kalman filter and an extended Kalman filter.
-Input-to-state stability analysis of the Jordan recurrent network-based estimator is discussed in Section II, while the method to find the corresponding Lyapunov functions is explained in Section II-B.
-The approach is then tested in Section IV on three different discrete-time systems, obtained from the discretization of ODEs, one linear and two nonlinear.
-This is followed by a discussion of the convergence and input-to-state stability of the results.
-II Input-to-state stability analysis
+We use this approach to establish a condition under which the error dynamics are stable if the original system is stable.
+Estimation of unstable systems is generally conducted in combination with a stabilizing feedback; our approach assumes that any necessary stabilization has been performed.
+The use of Jordan recurrent neural network for state estimation is described in Section II.
+Then in Section III an ISS approach is used to provide a characterization of a Lyapunov function for the error dynamics.
+Data generation and training are described in Section IV.
+The approach is illustrated in Section V with three different discrete-time systems.
+Figure 1: Jordan recurrent network (JRN) for state estimation showing output-to-hidden recurrent connections.
+II Jordan recurrent neural network for state estimation
 Consider a discrete-time nonlinear system
@@
 where $x^{(t)}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state vector at time-step $t$, $w^{(t+1)}\in\mathbb{R}^{n}$ is the process noise vector at time-step $t+1$, $y^{(t+1)}\in\mathcal{Y}\subseteq\mathbb{R}^{m}$ is the measurement vector at time-step $t+1$, $v^{(t+1)}\in\mathbb{R}^{m}$ is the measurement noise vector at time-step $t+1$, $x^{0}\in\mathbb{R}^{n}$ is the true initial condition and $\bar{x}^{0}\in\mathbb{R}^{n}$ is the initial condition noise.
-Here, $\mathcal{X}$ and $\mathcal{Y}$ are the state space and measurement space, respectively, and $n\in\mathbb{N}$ is the total number of states and $m\in\mathbb{N}$ is the number of states whose noisy measurements are available.
+Here, $\mathcal{X}$ and $\mathcal{Y}$ are the state space and measurement space, respectively, and $n\in\mathbb{N}$ is the total number of states and $m\in\mathbb{N}$ is the number of noisy measurements.
 The functions $f$ and $h$ are assumed to be sufficiently smooth.
-Our aim is to estimate the state vector $x^{(t+1)}$ by $\hat{x}^{(t+1)}$ based on the measurement vector $y^{(t+1)},$ previous state vector estimate $\hat{x}^{(t)}$ and the unknown noisy initial condition $x^{(0)}$.
+Our aim is to estimate the state vector $x^{(t+1)}$ by $\hat{x}^{(t+1)}$ based on the measurement vector $y^{(t+1)}$ and the previous state vector estimate $\hat{x}^{(t)}.$
 A JRN has recurrent connections from the output of the previous layer to the hidden state of the next layer.
@@
 Note that we are considering a bias-free network.
-To analyze the stability of the neural estimator, we first consider the case where there are no process noise and measurement noise in (1), i.e., $w^{(t+1)}=v^{(t+1)}=0$ for all $t$, and $\bar{x}^{0}=0$, as follows,
+III Input-to-state stability analysis
+To analyze the stability of the neural estimator, we first consider the case with no process and measurement noise in (1), i.e., $w^{(t+1)}=v^{(t+1)}=0$ for all $t$, and $\bar{x}^{0}=0$:
 $\begin{split}x^{(t+1)}&=f(x^{(t)}),\\ y^{(t+1)}&=h(x^{(t+1)}).\end{split}$ (3)
@@
 We introduce Lyapunov functions for the error dynamics, regarding $x^{(t)}$ as the input in (5).
-Definition II.1
+Definition III.1
 A continuous function $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{K}$-function if it is strictly increasing and $\alpha(0)=0$.
 It is a $\mathcal{K}_{\infty}$-function if it is a $\mathcal{K}$-function and $\alpha(r)\rightarrow\infty$ as $r\rightarrow\infty.$
-Definition II.2
-A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{KL}$-function if, for each fixed $t$, the function $\beta(\cdot,t)$ is a $\mathcal{K}$-function with respect to $r$, and for each fixed $r$, the function $\beta(r,\cdot)$ is decreasing with respect to $t$, and $\beta(r,t)\rightarrow 0$ as $t\rightarrow\infty.$
-Definition II.3
-The origin of (3) is globally asymptotically stable if, for each $\varepsilon>0$, there is a $\delta=\delta(\varepsilon)>0$ such that
+Definition III.2
+A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ is a $\mathcal{KL}$-function if, for each $t$, $\beta(\cdot,t)$ is a $\mathcal{K}$-function with respect to $r$, and for each $r$, $\beta(r,\cdot)$ is decreasing with respect to $t$, and $\beta(r,t)\rightarrow 0$ as $t\rightarrow\infty.$
+Definition III.3
+The origin is globally asymptotically stable for (3) if, for each $\varepsilon>0$, there is a $\delta=\delta(\varepsilon)>0$ such that
 $\|\xi\|<\delta\implies\|x^{(t)}\|<\varepsilon,\forall t\geq 0,\text{and}\lim_{t\rightarrow\infty}x^{(t)}=0.$ (6)
@@
 $\|x^{(t)}(\xi)\|\leq\beta(\|\xi\|,t)\quad\forall\xi\in\mathcal{X},\forall t\in\mathbf{Z}_{+}.$ (7)
-Definition II.4
-[24] The discrete-time system (5) is input-to-state stable (ISS) if there exists a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\gamma$, such that for each $t\in\mathbf{Z}_{+}$, the following inequality holds,
+Definition III.4
+[25] The discrete-time system (5) is input-to-state stable (ISS) if there exists a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\gamma$, such that for each $t\in\mathbf{Z}_{+}$,
 $|e^{(t)}(\eta,x)|\leq\beta(\|\eta\|,t)+\gamma(\|x\|),$ (8)
 for all $\eta\in\mathcal{E}$ and for all $x\in\mathcal{X}$.
-Definition II.5
+Definition III.5
 [25] A continuous function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$ is called an ISS-Lyapunov function for the discrete-time system (5) if there exist $\mathcal{K}_{\infty}$-functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a $\mathcal{K}$-function $\gamma$ such that
@@
 and
-$\begin{split}V(g(e^{(t)},x^{(t)}))-V(e^{(t)})&\leq-\alpha_{3}(\|e^{(t)}\|)+\gamma(\|x^{(t)}\|)\\ &\forall e^{(t)}\in\mathcal{E},\forall x^{(t)}\in\mathcal{X}.\end{split}$ (10)
-Theorem II.6
+$V(g(e^{(t)},x^{(t)}))-V(e^{(t)})$ $\leq-\alpha_{3}(\|e^{(t)}\|)+\gamma(\|x^{(t)}\|)$
+$\forall e^{(t)}\in\mathcal{E},\forall x^{(t)}\in\mathcal{X}.$ (10)
+Theorem III.6
 If the error system (5) with $x^{(t)}$ as input is ISS and the origin of the discrete-time system (3) is globally asymptotically stable, then the origin of the cascade system (3) and (5) is globally asymptotically stable.
@@
 The solutions of (3) and (5) satisfy:
-$\|x^{(t)}(\xi)\|\leq\beta_{1}(\|\xi\|,t),$ (11)
-$\|e^{(t)}(\eta,x)\|\leq\beta_{2}(\|\eta\|,t)+\gamma(\|x^{(t)}\|),$ (12)
+$\|x^{(t)}(\xi)\|$ $\leq\beta_{1}(\|\xi\|,t),$
+$\|e^{(t)}(\eta,x)\|$ $\leq\beta_{2}(\|\eta\|,t)+\gamma(\|x^{(t)}\|),$
 where $\beta_{1},\beta_{2}$ are $\mathcal{KL}$-functions. Thus,
@@
 Let $s^{(t)}$ denote the concatenation of the state $x^{(t)}$ and the error state $e^{(t)}$, and $\zeta$ denote the origin of this cascade system.
-We have $\|x^{(t)}\|\leq\|s^{(t)}\|$ and $\|e^{(t)}\|\leq\|s^{(t)}\|$, and $\|s^{(t)}\|\leq\|x^{(t)}\|+\|e^{(t)}\|$.
+We have $\|x^{(t)}\|\leq\|s^{(t)}\|$ and $\|e^{(t)}\|\leq\|s^{(t)}\|$, and $\|s^{(t)}\|\leq\|x^{(t)}\|+\|e^{(t)}\|$ (because for any 2 non-negative real numbers $a$ and $b,\sqrt{a^{2}+b^{2}}\leq a+b$).
 Defining $\beta(\cdot,\cdot)=\beta_{1}(\cdot,\cdot)+\beta_{2}(\cdot,\cdot)+\gamma(\beta_{1}(\cdot,\cdot))$, this yields
 $\|s^{(t)}(\zeta)\|\leq\beta(\|\zeta\|,t)\quad\forall t\in\mathbf{Z}_{+}.$ (12)
-It can be easily verified that $\beta$ is a $\mathcal{KL}$-function, which completes the proof. ∎
+It can be easily verified that $\beta$ is a $\mathcal{KL}$-function. ∎
 Two different approaches were used for finding the Lyapunov function, depending on whether the system is linear or nonlinear.
-II-A Linear systems
-Consider a linear discrete-time system defined by, for matrices $A$ and $H$,
+Linear Systems. Consider a linear discrete-time system defined by, for matrices $A$ and $H$,
 $\begin{split}x^{(t+1)}&=Ax^{(t)},\\ y^{(t+1)}&=Hx^{(t+1)}.\end{split}$ (13)
@@
 In this case, it is easy to show that both properties in Def.
-II.5 are satisfied with $\alpha_{1}(x)=\lambda_{min}(P)x^{2}$, $\alpha_{2}(x)=\lambda_{max}(P)x^{2}$, $\alpha_{3}(x)=\frac{1}{2}\lambda_{min}(Q)x^{2}$, and $\gamma(x)=(\frac{2|\mathcal{A}^{T}P\mathcal{B}|^{2}}{\lambda_{min}(Q)}+|\mathcal{B}^{T}P\mathcal{B}|^{2})x^{2}$  [24].
-II-B Nonlinear systems
-For nonlinear systems, a counterexample-guided method with verification provided by SMT solvers was used to synthesize the ISS-Lyapunov functions for the error system (5).
+III.5 are satisfied with $\alpha_{1}(x)=\lambda_{min}(P)x^{2}$, $\alpha_{2}(x)=\lambda_{max}(P)x^{2}$, $\alpha_{3}(x)=\frac{1}{2}\lambda_{min}(Q)x^{2}$, and $\gamma(x)=(\frac{2|\mathcal{A}^{T}P\mathcal{B}|^{2}}{\lambda_{min}(Q)}+|\mathcal{B}^{T}P\mathcal{B}|^{2})x^{2}$  [25].
+Nonlinear systems. For nonlinear systems, a counterexample-guided method with verification provided by SMT solvers was used to synthesize the ISS-Lyapunov functions for the error system (5).
 Inspired by [26], we use a one-hidden layer feed-forward neural network with zero bias terms for all layers to learn a Lyapunov function of the following form
@@
 We use the square function as the activation function $\sigma$ in this network, which results in sum-of-squares (SOS)-like quadratic Lyapunov functions.
-Consequently, we have $V(0)=0$.
+Then $V(0)=0$.
 As a result, a valid ISS Lyapunov function is attained when properties (9) and  (10) are satisfied.
@@
 Otherwise, it produces counterexamples that can be added to the training dataset of the neural network for finding Lyapunov function candidates.
-It is worth mentioning that due to the nature of the SMT solvers, we typically need to exclude a small region around the origin when verifying the falsification constraints.
-Around the origin, the linearization of the nonlinear system dominates, and we can implement the method for the linearized model to provide stability guarantees as detailed in Section II-A.
-We refer the readers to [26] for a detailed rigorous proof and algorithm.
+Due to the nature of the SMT solvers, we typically need to exclude a small region around the origin when verifying the falsification constraints.
+Around the origin, the linearization of the nonlinear system dominates, and we implement the method for the linearized model, described above, to provide stability guarantees.
+We refer the readers to [21] for a detailed proof and algorithm.
 The loss function is consistent with the falsification constraints:
-$\begin{split}L(\theta)&=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\max\bigg{(}0,V_{\theta}(g(e_{i},x_{j}))-V_{\theta}(e_{i})\\ &+\alpha_{3}(|e_{i}|)-\gamma(|x_{j}|)\bigg{)}+\max\bigg{(}0,V_{\theta}(e_{i})-\alpha_{2}(|e_{i}|)\bigg{)}\\ &+\max\bigg{(}0,\alpha_{1}(|e_{i}|)-V_{\theta}(e_{i})\bigg{)}.\end{split}$ (22)
-This is known as the positive penalty for the violation of conditions in Definition II.5.
-III Implementation
-The initial conditions are sampled over an interval.
-The initial condition is considered to be Gaussian.
-The covariance of the initial condition is $(0.01)\times I$ where $I$ is an identity matrix of appropriate dimensions.
-Similarly, process and measurement noises are assumed to be Gaussian with zero mean and covariance $0.01\times I$ where matrix $I$ is of appropriate dimensions.
-This is done for each sequence in the data set.
-Since the approach is state estimation, knowledge of the functions $f$ and $h$ in (1) is used in data generation.
-The data set is divided into three parts: training, validation and testing in the ratio $80:10:10$
-Figure 2: The graph represents a comparison between JRN and KF Error$(t),0\leq t\leq 100$ values of Jordan recurrent network based state estimation and the Kalman filter based state estimation for mass-spring damper system.
-The magnified graph within the graph provides a closer look at the error values from time $19$ to $20$ seconds.
-The x-axis of both the original and the magnified graphs represents time in seconds while the y-axis represents error at time $t.$ The blue dotted line graph represents the error values for JRN based estimates while the orange solid line represents the error values for the Kalman filter based estimates.
-III-A Training and validation
+$L(\theta)$ $=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}\max\bigg{(}0,V_{\theta}(g(e_{i},x_{j}))-V_{\theta}(e_{i})$
+$+\alpha_{3}(|e_{i}|)-\gamma(|x_{j}|)\bigg{)}+\max\bigg{(}0,V_{\theta}(e_{i})-\alpha_{2}(|e_{i}|)\bigg{)}$
+$+\max\bigg{(}0,\alpha_{1}(|e_{i}|)-V_{\theta}(e_{i})\bigg{)}.$ (20)
+This is known as the positive penalty for the violation of conditions in Definition III.5.
+IV Implementation
+The datasets are obtained by discretizing common ordinary differential equations.
+We use a zero order hold discretization for linear systems.
+For nonlinear continuous-time state space systems, we use the RK-45 discretization using Python’s $scipy.integrate.solve\_ivp$ function.
+The initial condition is considered to be Gaussian with mean sampled uniformly from the interval $[-1,1]\times[-1,1]\subset\mathbb{R}\times\mathbb{R}$ for each sequence and known covariance.
+Process and measurement noises are assumed to be Gaussian with zero mean and known covariance.
+For the systems under consideration, the above 3 covariance matrices are assumed to be $0.01\times I$ where $I$ is an identity matrix of appropriate dimensions.
+Also, knowledge of the functions $f$ and $h$ in (1) is used to generate data.
+For the linear system, we consider a total of $100$ sequences and $200$ for non-linear systems.
+The data set is divided into three parts: training, validation and testing in the ratio $80:10:10$ sequences.
+Each of the generated sequences is independent of the other sequences in the dataset.
 A custom network with forward propagation as in (2) is implemented.
 The weight matrix $W_{ay}$ is initialized using Xavier uniform distribution while weight matrices $W_{ya}$ and $W_{xa}$ are initialized to be (semi) orthogonal matrices using Pytorch’s $torch.nn.init$ module.
-The function $\sigma(z)$ is chosen to be equal to $z$ for linear systems and is chosen to be $\tanh(z)$ for nonlinear systems.
+The function $\sigma(z)$ is chosen to be equal to $z$ for the linear system and $\tanh(z)$ for the nonlinear systems.
 The backward propagation is implemented using PyTorch’s $backward()$ function.
-The loss function at time-step $t+1$ for each sequence is
+We consider the mean squared error loss function at time-step $t+1$ for each sequence
 $J_{i}^{(t+1)}(\phi)=\frac{1}{n}\sum_{i=1}^{n}{(x^{(t+1)}_{i}-\hat{x}^{(t+1)}_{i,\phi})}^{2},$ (21)
-where $1\leq i\leq n$ and $n$ is the number of states, $x^{(t+1)}_{i}$ represents the true value of state $x_{i}$ at time-step $t+1$ and $\hat{x}^{(t+1)}_{i,\phi}$ represents the estimated value of state $x_{i}$ at time-step $t+1$ by the JRN where $\phi$ denotes hyper-parameters of the network.
+where $1\leq i\leq n$ and $n$ is the number of states, $x^{(t+1)}_{i}$ represents the true value of state $x_{i}$ at time-step $t+1$ and $\hat{x}^{(t+1)}_{i,\phi}$ represents the estimated value of state $x_{i}$ at time-step $t+1$ by the JRN where $\phi$ denotes the weights of the network.
+Hyperparameter tuning was performed for all examples individually.
 Adam optimization is used for training the network.
-The optimal learning rate is chosen based on the network’s performance over a range of learning rates.
-The number of hidden units considered is $50$ and a batch size of $40$ is used for each example in Section IV.
+The optimal learning rate is chosen based on the network’s performance on validation dataset over a range of learning rates varying from $10^{-1}$ to $10^{-4}$.
+A batch size of $40$ and a hidden unit size of $50$ are considered for each example in Section V.
 Early stopping with a fixed patience value to decide the number of epochs needed is used.
 In each epoch, network’s performance on validation data is used as the measure to decide whether to proceed or keep training for another epoch.
-III-B Testing and comparison with KF and EKF
-We tested our estimators for 3 different systems in Section IV by comparing the results with KF for the linear system and EKF for the nonlinear systems.
-Standard implementations of KF and EKF are used, see for example [2].
-As mentioned earlier, $10\%$ of the data set is used for testing. We compare JRN and EKF graphically using average error at time $t$ over all features and all test sequences
+Maximum number of epochs is set to $600$ in case early stopping is not reached.
+For each lower validation loss value than the previous, the model is saved and after training reloaded to the one corresponding to the least validation loss.
+We test our estimators for 3 different systems in Section V.
+We compare the results with KF for the linear system as it is an optimal filter for systems with additive white Gaussian noise.
+For nonlinear systems with additive white Gaussian noise, the best filters are considered to be EKF and UKF.
+Hence, we compare our nonlinear systems results with them.
+Standard implementations of these 3 filters are used, see for example [1].
+For EKF, the equilibrium point around which the systems are linearized is considered to be the origin.
+We compare JRN, EKF and UKF graphically using average error at time-step $t$ over all features and all test sequences
 $\text{Error}(t)=\frac{1}{m_{test}n}\sum_{k=1}^{m_{test}}\sum_{i=1}^{n}(x_{i}^{(t)[k]}-\hat{x}_{i}^{(t)[k]})^{2},$ (22)
 where $m_{test}$ is the number of test sequences, $n$ is the number of states, $x_{i}^{(t)[k]}$ represents the true value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(t)[k]}$ represents the estimated value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence.
-We also use normalised mean square error (NMSE) to compare different methods.
+We also use root mean square error (RMSE) to compare different methods.
 It is calculated as
-$\text{NMSE}=\frac{1}{m_{test}nT}\sum_{k=1}^{m_{test}}\sum_{j=1}^{T}\sum_{i=1}^{n}(x_{i}^{(j)[k]}-\hat{x}_{i}^{(j)[k]})^{2},$ (25)
+$\text{RMSE}=\sqrt{\frac{1}{m_{test}nT}\sum_{k=1}^{m_{test}}\sum_{j=1}^{T}\sum_{i=1}^{n}(x_{i}^{(j)[k]}-\hat{x}_{i}^{(j)[k]})^{2}},$ (23)
 where $m_{test}$ is the number of test sequences, $T$ is the number of time-steps in each sequence, $n$ is the number of states, $x_{i}^{(j)[k]}$ represents the true value of state $x_{i}$ at time-step $j$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(j)[k]}$ represents the estimated value of state $x_{i}$ at time-step $j$ for the $k^{th}$ sequence.
-IV Numerical results
-In this section, we illustrate the proposed method with several examples.
-All three discrete-time systems were obtained by discretizing common ordinary differential equation examples.
-We use a zero order hold discretization for linear systems.
-For nonlinear continuous-time state space systems, we use the RK-45 discretization using Python’s $scipy.integrate.solve\_ivp$ function.
-A remote Windows server with T4 GPU, courtesy of math faculty computing facility (MFCF), University of Waterloo, was used for all numerical simulations.
-(a) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 50$ seconds for down pendulum.
-(b) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 20$ seconds for reversed Van der Pol oscillator.
-Figure 3: The graphs represent a comparison of Error$(t)$ values of Jordan recurrent network-based state estimation and extended Kalman filter-based state estimation for nonlinear dynamical systems.
-The magnified graph within each graph provides a closer look at the error values from time $19$ to $20$ seconds.
-The x-axis of both the original and the magnified graphs represents time in seconds, while the y-axis represents error at time $t$.
-The blue dotted line graph represents the error values for JRN-based estimates, while the orange solid line represents the error values for extended Kalman filter-based estimates.
-IV-A Mass-Spring-Damper system
-Consider a zero order hold discretization with $\Delta t=0.1s$ of the mass-spring damper system
-$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{k}{m}x_{1}(t)-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (26)
-where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity respectively of the spring.
-The parameters are $m=10kg,$ $b=6kg/sec$ and $k=800kg/{sec}^{2}.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
-We consider $100$ sequences from 00 to $100$ seconds each.
-The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ Early stopping was implemented with a patience of $5$.
-We use Adam optimization for training with a learning rate of $10^{-4}.$ There were $4216$ epochs for training.
-Fig.
-2 shows the error at each time-step averaged all test sequences for JRN and EKF.
-As can be seen from Fig.
-2, the error for JRN converges to 00 as fast as for the KF.
-Also, the NMSE for JRN is $0.0631$ and for KF is $0.0653$ which is very close.
-Recall that KF is the optimal estimator for this system.
-Since the initial condition is not assumed to be known by the estimator, there is a spike in the initial values of the error for the KF and the JRN but the spike in KF error is considerably larger than that of JRN.
-The magnified graph in Fig.
-2 shows that after convergence JRN error is very close to that the KF error.
-For this 2-dimensional linear system, the origin is asymptotically stable.
-We then implement the method in II-A for linear systems.
-Setting $Q=I$, it is easy to establish that $P=\begin{bmatrix}1.0236&-0.0934\\ -0.0934&69.4428\end{bmatrix}$ With this ISS-Lyapunov function, by Theorem II.6, both the original system and the error system are asymptotically stable.
+Figure 2: Figures 2(a), 2(b) and 2(c) show the training and validation values of JRNs for the mass spring damper system, down pendulum and reversed Van der Pol oscillator respectively.
+Figures 2(d) 2(e) and 2(f) show average errors at each time-step for JRN, EKF and UKF state estimates for mass spring damper system, down pendulum and reversed Van der Pol oscillator respectively.
+Figure 3: Learned ISS Lyapunov function for a down pendulum.
+TABLE I: Training hyperparameters
+State space system Time-steps Learning rate Patience
+Mass spring 300 0.01 10
+Down pendulum 300 0.01 25
+r. Van der Pol 200 0.001 25
+TABLE II: Parameters
+State space system Parameters
+Mass spring $m=10$ kg, $b=6$ kg/sec, $k=800$ kg/sec2
+Down pendulum $m=2$ kg, $b=0.9$ kg/sec, $l=1m.$
+V Examples
+In this section, we illustrate the proposed method with 3 discrete-time systems obtained as explained in IV.
+The ODEs considered are mass-spring damper system, down pendulum and reversed Van der Pol oscillator.
+The measurement vector at time-step $t$ for each of the systems is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $x_{1}(t)$ refers to the position and $\nu^{(t)}$ is the measurement noise at time-step $t$.
+Important parameters for each of these systems are summarised in Tables I and II.A remote Ubuntu server was used for all calculations.
+The codes used are available on Github at https://github.com/avneetkaur96/JRNs-for-state-estimation.
+TABLE III: Root mean square error (RMSE) for E(KF), UKF and JRN
+State space Estimator
+system E(KF) UKF JRN
+Mass spring 0.5132 - 0.5268
+Down pendulum 0.1594 0.1693 0.1298
+r. Van der Pol 0.1153 0.1158 0.1102
+TABLE IV: Best epoch and training time for JRN
+State space system Best Epoch Training time(in sec)
+Mass spring 135 20
+Down pendulum 313 101
+r. Van der Pol 96 32
+The performance of the JRN and the KF for mass-spring system state estimation is summarized in Figures 2(d) and 2(a).
+As shown in Figure 2(d), both the JRN and KF achieve rapid error reduction within the first few time steps, with the KF maintaining slightly lower error overall as expected due to its optimal nature.
+The JRN closely tracks the KF’s performance after the initial steps, indicating effective learning of the system dynamics.
+The difference of estimation in the first few timesteps can be attributed to the fact that the KF is provided with an estimate of the initial condition while the JRN has no information of the initial condition.
+The training and validation loss curves for the JRN (Figure 2(a)) demonstrate fast convergence, with loss stabilizing after approximately 40 epochs.
+The close alignment of training and validation loss throughout training indicates good generalization and minimal overfitting.
+The RMSE on the test set is $0.5132$ for the KF and $0.5268$ for the JRN (Table III, confirming that the JRN provides comparable estimation accuracy.
+The model was restored to its state at the best epoch for testing.
+For this system, the origin is asymptotically stable.
+We then implement the method for stability of linear systems.
+Setting $Q=I$, it is easy to establish that $P=\begin{bmatrix}45.504&-3.5737\\ -3.5737&1.4461\end{bmatrix}.$ With this ISS-Lyapunov function, by Theorem III.6, both the original system and the error system are asymptotically stable.
 This proves the stability of the NN estimator error.
-IV-B Down pendulum
-Consider an RK-45 discretization with $\Delta t=0.01s$ of the down pendulum
-$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{g}{l}\sin({x_{1}(t)})-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (27)
-where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity of the pendulum respectively.
-The parameters are $g=9.8m/{sec}^{2},$ $m=2kg,$ $b=0.9kg/sec$ and $l=1m.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
-We consider $200$ sequences from 00 to $50$ seconds each.
-The mean of the initial condition is chosen randomly from $[-2,2]\times[-2,2].$ Early stopping was implemented with a patience of $50$.
-The validation loss reached early stopping patience at $2984$ epochs.
-We further used Adam optimization for training with a learning rate of $10^{-4}.$
-The error at each time-step, averaged over all features and all test sequences, is shown for JRN and EKF in Fig.
-3(a).
-The error for JRN converges to 00 as fast as that for the EKF.
-Also, the NMSE for JRN is $0.0627$ while for EKF it is $0.0647$ The magnified graph in Fig.
-3(a) shows that after convergence JRN error and EKF error are quite close.
-The initial spike in JRN error is not as high as the EKF error.
-The error becomes very small as $t$ goes to $\infty$, despite the presence of noise.
-For this nonlinear system, an ISS Lyapunov function was learned using a neural network, as shown in Fig.
-4, which is verified by an SMT solver, dReal [27], on a $[-4,4]$ for the error state.
+As shown in Figures 2(e) and 2(f), all three approaches exhibit a rapid reduction in error within the first 25 time-steps, stabilizing thereafter.JRN achieves the lowest estimation error over time, outperforming both EKF and UKF for the down pendulum.
+The errors for both systems and for all methods stabilize after approximately $50$ time-steps, with JRN maintaining a consistently lower error throughout for the down pendulum.
+Figures 2(b) and 2(c) demonstrate the convergence of the JRN training process.
+Both training and validation losses decrease steadily, indicating good generalization and absence of overfitting.
+The JRN achieves the lowest RMSE compared to the EKF and UKF for both the nonlinear systems, indicating improved accuracy (See Table III).
+These results highlight the effectiveness and efficiency of the JRN in nonlinear state estimation tasks.
+For both nonlinear systems, an ISS Lyapunov function was learned using a neural network and verified by an SMT solver.
+For the down pendulum, this function is shown in Fig.
+3, verified by an SMT solver, dReal [27], on a $[-2,2]\times[-2,2]$ for the error states.
 In this case, we used $\alpha_{1}(\cdot)=\alpha_{3}(\cdot)=0.01|\cdot|$ and $\alpha_{2}(\cdot)=\gamma(\cdot)=100|\cdot|$.
-Figure 4: The learned ISS Lyapunov function for the down pendulum example.
-IV-C Reversed Van der Pol oscillator
-Consider an RK-45 discretization with $\Delta t=0.1s$ of the reversed Van der Pol oscillator
-$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}-x_{2}(t)\\ x_{1}(t)+(({x_{1}}(t))^{2}-1)x_{2}(t)\end{bmatrix}$ (28)
-where $x_{1}(t)$ and $x_{2}(t)$ is the position and velocity of the oscillator.
-The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
-We consider $200$ sequences from 00 to $20$ seconds each.
-The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ The training was done for $35000$ epochs as early stopping patience of $7000$ was not achieved.
-The validation error continued to decrease.
-Adam optimization is used for training with a learning rate of $10^{-4}.$ As can be seen from Fig.
-3(b), the JRN error converges to zero much faster than the EKF error.
-Also, the NMSE for JRN is $0.0484$ while for EKF is $0.1025$ This is considerable difference.
-The magnified graph shows that even after convergence, the JRN error is better than EKF error.
-This example suggests that when a network is trained for a larger number of epochs, the JRN error can converge much faster than the EKF error.
-For this system, in a similar manner to the previous example, we synthesized an ISS-Lyapunov function on $[-2,2]$ for the error states, which provides the asymptotic stability for the error term by Theorem II.6.
-V Conclusions and future work
-In this paper, we demonstrated that when knowledge of functions $f$ and $h$ in (1) is available or the data of the real systems can be properly sampled, state estimation can be performed by using Jordan recurrent networks.
-The Jordan recurrent network based estimator helped us achieve results as good as KF and EKF.
-We also showed the stability of the error dynamics of the neural estimator using an ISS approach, by taking advantage of the structural properties of an unbiased JRN.
+This provides the asymptotic stability for the error dynamics by Theorem III.6.
+VI Conclusions and future work
+The Jordan recurrent network based estimator achieved results as good as KF for the linear system and better than EKF and UKF for the chosen nonlinear systems.
+While the training time is a drawback in case of neural networks, the testing time of JRN-based estimators is much lower than the classical approaches of UKF and EKF.
+Current work focuses on extending the Jordan architecture to a long short-term neural network [14].
+The main contribution of this paper is to verify the stability of the error dynamics of the RNN-based estimator using an ISS approach, by taking advantage of the structural properties of an unbiased JRN and recent results in learning Lyapunov functions.
 For linear systems, quadratic Lyapunov equations can be computed analytically for the error dynamics to provide stability.
 For nonlinear systems, a counter-example guided method was used to learn an ISS Lyapunov function.
-Furthermore, the effectiveness of the proposed method has been illustrated with three 2-dimensional systems.
-We will extend our approach to higher dimensional systems in future research.
-Another potential future research direction is to consider stability for the case where biases are non-zero.
-Additional further work is to extend the results on the convergence of discrete-time recurrent neural networks as shown in [28] to JRNs.
+The linear approach is used near the equilibrium.
+The effectiveness of the method has been illustrated with three examples.
+Future work will extend this approach to higher-order systems.
+Addressing the issue of estimation of unstable systems is also of interest.
+Another potential future research direction is to consider stability for the case where biases are non-zero, and for networks with long short term memory.